OpenStudy (anonymous):

(x^2-x-9)^3/4-13=14 solve the equation with rational exponents

myininaya (myininaya):

\[(x^2-x-9)^\frac{3}{4}=27\] \[(x^2-x-9)=27^\frac{4}{3}\] \[(x^2-x-9)=3^4\] \[x^2-x-9=81\] \[x^2-x-90=0\] \[(x-10)(x+9)=0\]

myininaya (myininaya):

\[x=10 or x=-9\]

OpenStudy (anonymous):

add 13 to both sides is a start. get \[(x^2-x-9)^{\frac{3}{4}}=27\] take the cube root get \[\sqrt[4]{(x^2-x-9)}=3\] raise to the power of 4 get \[x^2-x-9=81\] then solve

OpenStudy (anonymous):

or be quick like myininaya (show off!)

myininaya (myininaya):

check both since you have an even root

myininaya (myininaya):

obviously 10 is god now check -9

myininaya (myininaya):

and -9 is good

OpenStudy (anonymous):

aren't you tired after a long day of calc IX?

myininaya (myininaya):

i'm suppose to be reading something

OpenStudy (anonymous):

torturing your poor dears by making the find derivatives using triangles? you cover series in that class?

myininaya (myininaya):

no series is cover in our cal 3 class i need to review that stuff its been forever since i talked about convergence of a series

OpenStudy (anonymous):

too bad they took the one fun thing out and left those stultifyingly boring "techniques of integration" in entire content of class on the back cover of stewart

myininaya (myininaya):

oh yeah i had a nap

OpenStudy (anonymous):

zzzzzzzzzzzzzzz